OpenStudy (anonymous):

Can someone please help me with my prelab questions. Write a protocol for how to prepare 100mL of a 0.50M AgNO3 solution

OpenStudy (chmvijay):

can u tell me AgNO3 molar mass

OpenStudy (anonymous):

Its 169.875g/mol

OpenStudy (abb0t):

The solution contains 0.1 liter \(\times\) 0.5M = _______ mol AgNO\(_3\). Now, calculate the mass: _______ mol \(\times\) 169.873 \(\frac{g}{mol}\) = _____ g So dilute _____ g AgNO\(_3\) in 100 ml water.

OpenStudy (anonymous):

9g of AgNO3

OpenStudy (chmvijay):

169.873 -1000ml -1M 16.9873-100l-1M 8.5----100ml-0.5M

OpenStudy (anonymous):

Thanks

OpenStudy (chmvijay):

yw :)